# Social Optimization in Noncooperative Games under Central Regulation

### v1

@@ II-A Problem Formulation
$\mathcal{G}$ is strongly connected. The adjacency matrix $\mathcal{A}$ is doubly stochastic, i.e., $\mathcal{A}\bm{1}_{N}=\mathcal{A}^{\textit{T}}\bm{1}_{N}=\bm{1}_{N}$.
Under Assumption 3, there exists $\bar{\sigma}\in(0,1)$ such that $\sigma_{N-1}(\mathcal{A})\leq\bar{\sigma}$.
Generally speaking, by virtue of the properties of objective functions and constraint sets of the players, various methods can be employed to compute NEs of the low-level game (1). Hence, our main attention shifts towards addressing the high-level optimization problem (3). From the preceding analysis, it is clear that to minimize the high-level function $f(x(\theta),\theta)$ with respect to $\theta$, one needs to firstly solve the low-level game problem (1), while its corresponding optimal decision $x(\theta)$ in turn depends on $\theta$. This mutual coupling between $x$ and $\theta$ renders problem (II-A) intrinsically hard to solve. Naturally, a direct and efficient approach involves gradient-based methods, where an implicit gradient of the high-level function $f(x(\theta),\theta)$ is required to be computed by utilizing the chain rule, as follows
@@
Subsequently, the high-level decision variable $\theta$ is updated utilizing the projected gradient descent method. However, several challenges arise:
@@
    For a particular $\theta\in\mathbb{R}^{n}$, obtaining a closed-form characterization of $x(\theta)$ is often computationally expensive or even unavailable unless the functions $f_{i}$, $i\in[N]$ take very special forms. Hence, this typically limits the applicability of standard first-order and zeroth-order schemes.
@@
    The computation of implicit gradient $\nabla f(\theta,x(\theta))$ requires not only access to $x(\theta)$ but also the assumption of differentiability of $x(\theta)$. Unfortunately, as demonstrated by Example 1 below, $x(\theta)$ and hence $f(x(\theta),\theta)$ are non-differentiable in general, rendering the gradient-based methods infeasible. Furthermore, one may not compute subgradients or Clarke generalized gradients easily in such settings either [30]. As a consequence, the design and analysis of computational methods for addressing problem (II-A) are considerably challenging.
@@
    The function $f(x(\theta),\theta)$ may exhibit nonconvexity in $\theta$, as demonstrated by Example 1 below. This imposes significant challenges in the convergence analysis and, especially, in deriving the convergence rate.
Example 1.
Consider a noncooperative game with $2$ players. The cost function of player $i$ is $f_{i}(x_{i},x_{-i},\theta)=x_{i}^{2}-2x_{-i}-2x_{i}\theta$ and the strategy set is $X_{i}=[2/3,1]$, $\forall i\in[2]$. The feasible set of the central regulator is $\Theta=[0,1]$. Based on the optimal condition (11), it can be derived that
Through simple calculation, one has
$f(x(\theta),\theta)$ $=\sum_{i=1}^{2}f_{i}(x(\theta),\theta)=\left\{\begin{aligned} &-\frac{4}{3}\theta-\frac{8}{9},&\text{if}\ \theta\in[0,2/3],\\ &-2\theta^{2}-4\theta,&\text{if}\ \theta\in[2/3,1].\end{aligned}\right.$
II-B Randomized Smoothness
@@
where $\xi>0$ is the smoothing parameter. Then a two-point sampling gradient estimator is proposed as [31]
^$\hat{\nabla}g(z)=\frac{n}{\xi}(g(z+\xi u)-g(z))u,$ (14)
@@
The following lemma shows some properties of $\hat{g}$ and ^$\hat{\nabla}g$.
@@
    $\hat{g}(z)$ is differentiable even when $g$ is not, it holds that
@@
    ^$\|\hat{\nabla}g(z)\|\leq nL_{g}.$ (16)
II-C Moreau Smoothness
@@
The Moreau smoothing $\hat{\chi}_{\Theta}$ of $\chi_{\Theta}$ is defined as [32]
@@
where $\xi$ is the smoothing parameter. By Theorem 6.60 in [32], $\hat{\chi}_{\Theta}$ is differentiable and
@@
Intuitively, the randomized smoothed version $\hat{g}$ and Moreau smoothed variant $\hat{\chi}_{\Theta}$, as substitutes for original functions $g$ and $\chi_{\Theta}$, respectively, are typically introduced to contend with the nonsmoothness, which greatly help develop a zeroth-order algorithm.
III MAIN RESULTS
In this section, a zeroth-order algorithm for addressing the constrained, nonsmooth, and nonconvex problem (II-A) is developed, which allows for utilizing inexact solutions of the low-level game. Moreover, the convergence rate of the proposed algorithm are rigorously derived.
For notational convenience, define $F_{i}(\theta):=f_{i}(x(\theta),\theta))$ and $F(\theta):=\sum_{i=1}^{N}F_{i}(\theta)$. Note that the constrained optimization problem (3) is equivalent to
$\min_{\theta\in\mathbb{R}^{n}}\mathbf{F}(\theta):=F(\theta)+\chi_{\Theta}(\theta).$ (18)
Consider a smoothed unconstrained approximation of (3) based on (13) and (17), given as
@@ III MAIN RESULTS
Before proceeding, we formalize the relationship between the original nonsmooth problem (II-A) and its smoothed counterpart (19). Toward this end, the $\delta$-Clarke generalized gradient of $\mathbf{F}$ at $\theta$ is introduced [33], denoted by $\partial_{\delta}\mathbf{F}(\theta)$, as follows
$\partial_{\delta}\mathbf{F}(\theta):=\operatorname{conv}\{\zeta\ |\ \zeta\in\partial\mathbf{F}(\theta^{\prime}),\|\theta^{\prime}-\theta\|\leq\delta\}.$ (20)
Basically, the set $\partial_{\delta}\mathbf{F}(\theta)$ is shown to be nonempty, compact, and convex [33]. Inspired by Proposition 2 in [30], the following lemma is shown.
@@
Consider problem (II-A) and let Assumptions 1 and 2 be satisfied. Then for any $\delta>0$, if $\nabla\hat{\mathbf{F}}(\theta)=\bm{0}$ and $\xi\leq\frac{\delta}{\max\{2,nNL_{F}\}}$, there hold $\bm{0}\in\partial_{\delta}\mathbf{F}(\theta)$ and $\|\theta-\Pi_{\Theta}(\theta)\|\leq\xi nNL_{F}$.
@@
In order to deal with the smoothed problem (19), by virtue of Lemma 2, for every $i\in[N]$, two point sampling gradient approximation of $\hat{F}_{i}(\theta)$ is
^$\hat{\nabla}F_{i}(\theta):=\frac{n}{\xi}(f_{i}(x(\theta+\xi u),\theta+\xi u)-f_{i}(x(\theta),\theta))u.$
Hence
^$\hat{\nabla}F(\theta)=\frac{n}{\xi}\sum_{i=1}^{N}(f_{i}(x(\theta+\xi u),\theta+\xi u)-f_{i}(x(\theta),\theta))u.$ (21)
Based on the preceding discussion, it is impractical to obtain the exact solution $x(\theta)$. We carefully address this challenge by introducing inexact evaluation of $x(\theta)$, where players are allowed to compute an $\varepsilon$-approximation solution $x_{\varepsilon}(\theta)$ satisfying
@@
^$\hat{\nabla}_{\varepsilon}F_{i}(\theta)=\frac{n}{\xi}(f_{i}(x_{\varepsilon}(\theta+\xi u),\theta+\xi u)-f_{i}(x_{\varepsilon}(\theta),\theta))u.$
@@
4      Apply Algorithm 2 to receive inexact evaluations $x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k})$ and $x_{\varepsilon_{k}}(\theta_{k})$.
5      Evaluate the inexact zeroth-order gradient approximation ^$\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ by (23).
@@
In a nutshell, an inexact smoothing-enabled zeroth-order method is developed in Algorithm 1, while an inexact solution of $x(\theta)$ is computed by a fully distributed NE seeking algorithm, outlined in Algorithm 2, which was inspired by [7]. At each iteration $k$, firstly, the central regulator select vector $u_{k}\in\mathbb{S}^{n}$ independently and uniformly at random and then send $\theta_{k}+\xi u_{k}$ and $\theta_{k}$ to all players. Simultaneously, the central regulator makes an inexact call to Algorithm 2 twice to compute $x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k})$ and $x_{\varepsilon_{k}}(\theta_{k})$. Specifically, at time $t$, each player $i$ maintains variables $x_{i,t}$ and ${\bf x}_{i,t}^{j}$ to represent its own decision and estimate of player $j$’s decision, respectively. Define ${\bf x}_{i,t}^{i}:=x_{i,t}$ and also ${\bf x}_{j,t}^{-i}:=\operatorname*{col}\{{\bf x}_{j,t}^{l}\}_{l\in[N]\setminus i}$. Let ${\bf x}_{i,t}:=\operatorname*{col}\{{\bf x}_{i,t}^{j}\}_{j\in[N]}\in\mathbb{R}^{d}$ and ${\bf x}_{t}:=\operatorname*{col}\{{\bf x}_{i}\}_{i\in[N]}\in\mathbb{R}^{Nd}$. In the partial-decision information setting, players update their estimates based on the information received from their neighbors by applying dynamic consensus technique (25a), and subsequently adjust their own decisions through a gradient descent step (25b). Herein, it is mentioned that the partial gradient $\nabla_{i}f_{i}$ is evaluated on the local estimate $\hat{{\bf x}}_{i,t}$, while not employing the actual decision $x_{i,t}$. After receiving the estimated feedback $x_{\varepsilon_{k}}(\tilde{\theta}_{k}):=\operatorname*{col}\{x_{i,t_{k}}\}_{i\in[N]}$ from the players, the inexact zeroth-order gradient ^$\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ is computed as
^$\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})=\sum_{i=1}^{N}\frac{n}{\xi}(f_{i}(x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k}),\theta_{k}+\xi u_{k})-f_{i}(x_{\varepsilon_{k}}(\theta_{k}),\theta_{k}))u_{k}.$ (23)
At last, the central regulator updates its decision using gradient descent
where $\alpha>0$ is the step size to be determined.
@@
2      for $i=1,...,N$ in parallel do
3           Receive $\tilde{\theta}_{k}$ (that is, $\theta_{k}+\xi u_{k}$ or $\theta_{k}$) from the central regulator.
@@ 4           Update
Players send $x_{i,t_{k}},i\in[N]$ to the central regulator.
Algorithm 2 Fully Distributed NE Seeking
@@
Utilizing inexact estimation of the NE of the low-level game inherently introduces a bias in the approximation of the zeroth-order gradient. This bias, if uncontrolled, may undesirably propagate and impact the overall performance of the algorithm. Notably, the degree of inexactness in $x_{\varepsilon_{k}}(\tilde{\theta}_{k})$ is precisely controlled by the termination criterion $t_{k}$ in Algorithm 2. This criterion will be formally derived in subsequent Theorem 1 to establish the convergence rate. Additionally, note that the central regulator in the high-level does not need to acquire all information on the players. In fact, at moment $k+1$, only two function values $f_{i}(x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k}),\theta_{k}+\xi u_{k})$ and $f_{i}(x_{\varepsilon_{k}}(\theta_{k}),\theta_{k})$ are required to be disclosed by the central regulator to update its decision $\theta_{k}$. In addition, it is noteworthy that in the setting of partial-decision information regarding the underlying game, each individual strategy and gradient of objective function are not accessible to other players. Consequently, every player must exert extra effort to estimate strategies of other players and hence steps (25a) and (25c) are introduced in Algorithm 2. However, during the practical implementation of Algorithm 1, it is only necessary to employ simple gradient descent method to update the players’ decision variables to obtain approximate solution $x_{\varepsilon_{k}}(\tilde{\theta}_{k})$ of the low-level game, which will simplify Algorithm 1 greatly.
@@
[7, Lemma 1] Under Assumptions 1-iii) and 2-i), for any $\theta\in\mathbb{R}^{n}$, the mapping $\mathbf{G}$ is $l^{\prime}$-Lipschitz continuous with respect to ${\bf x}$ for some $\mu\leq l^{\prime}\leq l$, that is,
@@
To proceed further analysis, the convergence result of Algorithm 2 for resolving the low-level game (1) is presented. The proof can be obtained following Theorem 1 in [7].
@@
with $a:=\bar{\sigma}^{2}(2ll^{\prime}+l^{\prime 2}+4\mu l^{\prime}+2l^{2})+2(l^{2}l^{\prime 2}+\mu l^{\prime 2}+2l^{2}l^{\prime 2})\bar{\sigma}^{2}+2l^{2}l^{\prime 2}\bar{\sigma}^{2}$, then $x_{t}$ converges to $x(\tilde{\theta}_{k})$ at a linear rate. Specifically,
@@
Define the natural filtration
$\mathscr{F}_{0}:=\{\emptyset,\Omega\},\ \mathscr{F}_{k}:=\sigma\{\theta_{0},u_{t}:0\leq t\leq k-1\},\ \forall k\geq 1.$
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]$ $\leq\frac{4(\hat{\mathbf{F}}(\theta_{0})-\hat{\mathbf{F}}^{*})}{\alpha K}+8N^{2}n^{2}L_{F}^{2}\alpha\frac{nNL_{F}+1}{\xi}$

### v2

@@ 2.1 Problem formulation
$\mathcal{G}$ is strongly connected. The adjacency matrix $\mathcal{A}$ is doubly stochastic, i.e., $\mathcal{A}\bm{1}_{N}=\mathcal{A}^{\top}\bm{1}_{N}=\bm{1}_{N}$.
Assumption 3 is a necessary condition for ensuring that players can estimate all other players’ strategies. Under Assumption 3, there exists $\bar{\sigma}\in(0,1)$ such that $\sigma_{N-1}(\mathcal{A})\leq\bar{\sigma}$.
Generally speaking, by virtue of the properties of objective functions and constraint sets of the players, various methods can be employed to compute the NE of the low-level game (1). Hence, our main focus shifts to addressing the high-level optimization problem in (2.1). From the preceding analysis, it is apparent that to minimize the high-level function $f(x(\theta),\theta)$ with respect to $\theta$, where $x(\theta)$ is the NE of the low-level game (1), one needs to firstly solve the low-level game problem (1), while its corresponding optimal decision $x(\theta)$ in turn depends on $\theta$. This mutual coupling between $x$ and $\theta$ renders problem (2.1) intrinsically hard to solve. Naturally, a direct and efficient approach involves gradient-based methods, in which an implicit gradient of the high-level function $f(x(\theta),\theta)$ is required to be computed by utilizing the chain rule, as follows
@@
However, several challenges arise:
@@
    For a particular $\theta\in\mathbb{R}^{n}$, obtaining a closed-form characterization of $x(\theta)$, where $x(\theta)$ is the NE of the low-level game (1), is often computationally expensive or even unavailable unless the functions $f_{i}$, $i\in[N]$ take a very special form. Consequently, this typically limits the applicability of standard first-order and zeroth-order schemes.
@@
    The computation of implicit gradient $\nabla f(\theta,x(\theta))$ requires not only access to $x(\theta)$ but also the assumption of differentiability of $x(\theta)$. Unfortunately, as demonstrated by Example 1, $x(\theta)$ and hence $f(x(\theta),\theta)$ are non-differentiable in general, rendering the gradient-based methods infeasible. As a result, the design of computational methods for addressing problem (2.1) remains challenging.
@@
    The function $f(x(\theta),\theta)$ may exhibit nonconvexity in $\theta$, as demonstrated by Example 1. This, in turn, poses another challenge in developing an effective algorithm with convergence guarantees, especially in establishing the convergence rate.
2.2 Randomized smoothness
@@
where $\xi>0$ is the smoothing parameter. Then a two-point sampling gradient estimator is proposed as [yi2021online]
$\hat{\nabla}g(z)=\frac{n}{\xi}(g(z+\xi u)-g(z))u,$ (13)
@@
The following lemma shows some properties of $\hat{g}$ and $\hat{\nabla}g$.
@@
    $\hat{g}(z)$ is differentiable even when $g$ is not. Especially, it holds that
@@
    $\|\hat{\nabla}g(z)\|\leq nL_{g}.$ (15)
2.3 Moreau Smoothness
@@
The Moreau smoothing $\hat{\chi}_{\Theta}$ of $\chi_{\Theta}$ is defined as [Beck2017]
@@
where $\xi$ is the smoothing parameter. By Theorem 6.60 in [Beck2017], $\hat{\chi}_{\Theta}$ is differentiable and
@@
Intuitively, the randomized smoothed version $\hat{g}$ and Moreau smoothed variant $\hat{\chi}_{\Theta}$, as substitutes for original functions $g$ and $\chi_{\Theta}$, respectively, are typically introduced to contend with the nonsmoothness, which greatly help to develop a zeroth-order algorithm.
3 Main results
In this section, a zeroth-order algorithm for addressing the constrained, nonsmooth, and nonconvex problem (2.1) is developed, which allows for leveraging inexact solutions of the low-level game. Moreover, the convergence rate of the proposed algorithm is rigorously derived.
For notational convenience, define
$F_{i}(\theta):=f_{i}(x(\theta),\theta)),\ F(\theta):=\sum_{i=1}^{N}F_{i}(\theta),$
where $x(\theta)$ is the NE of the low-level game (1). To proceed, we first note that the constrained optimization problem (2.1) is equivalent to
$\min_{\theta\in\mathbb{R}^{n}}\mathbf{F}(\theta):=F(\theta)+\chi_{\Theta}(\theta),$ (17)
where
Consider a smoothed approximation of (17) based on (12) and (16), given as
@@ 3 Main results
For nonsmooth and nonconvex optimization problem (17), it is shown in [zhang2020complexity] that no algorithm could find an $\epsilon$-stationary point, i.e., a point $\theta$ for which $\min\{\|\zeta\|\mid\zeta\in\partial\mathbf{F}(\theta)\}\leq\epsilon$ in finite time. Hence, as a tractable and reasonable optimality criterion, a relaxed notion ($\delta,\epsilon$)-stationarity is considered for a vector $\theta$ satisfying $\min\{\|\zeta\|\mid\zeta\in\partial_{\delta}\mathbf{F}(\theta)\}\leq\epsilon$, where the set $\partial_{\delta}\mathbf{F}(\theta):=\operatorname{conv}\{\zeta\ |\ \zeta\in\partial\mathbf{F}(\theta^{\prime}),\|\theta^{\prime}-\theta\|\leq\delta\}$ denotes the $\delta$-Clarke generalized gradient of $\mathbf{F}$ at $\theta$ [gold1977].
Based on the analysis above, the goal of this paper is to develop a provably convergent algorithm to find a ($\delta,\epsilon$)-stationary point of function $\mathbf{F}$. Before proceeding, we formalize the relationship between the original nonsmooth problem (2.1) and its smoothed counterpart (18).
@@
Consider problem (2.1) and let Assumptions 1 and 2 be satisfied.
  • i)
    If $\Theta=\mathbb{R}^{n}$, then for any $\xi>0$, it has $\nabla\hat{F}(\theta)\in\partial_{2\xi}F(\theta)$.
  • ii)
    For any $\delta>0$, if $\nabla\hat{\mathbf{F}}(\theta)=\bm{0}$ and $\xi\leq\frac{\delta}{\max\{2,nNL_{F}\}}$, then it holds that $\bm{0}\in\partial_{\delta}\mathbf{F}(\theta)$ and $\|\theta-\Pi_{\Theta}(\theta)\|\leq\xi nNL_{F}$.
Proof.
See Appendix A. ∎
@@
Next, let us proceed to the algorithm design. In order to deal with the smoothed problem (18), by virtue of Lemma 1, for every $i\in[N]$, a two point sampling gradient approximation of $\hat{F}_{i}(\theta)$ is
$\hat{\nabla}F_{i}(\theta)=\frac{n}{\xi}(f_{i}(x(\theta+\xi u),\theta+\xi u)-f_{i}(x(\theta),\theta))u.$
Hence,
$\hat{\nabla}F(\theta)$ $=\sum_{i=1}^{N}\hat{\nabla}F_{i}(\theta)$
$=\sum_{i=1}^{N}\frac{n}{\xi}(f_{i}(x(\theta+\xi u),\theta+\xi u)-f_{i}(x(\theta),\theta))u.$ (19)
Based on the preceding discussion, it is impractical to obtain the exact solution $x(\theta)$, where $x(\theta)$ is the NE of the low-level game (1). We carefully address this challenge by introducing inexact evaluation of $x(\theta)$, where players are allowed to compute an $\varepsilon$-approximation solution $x_{\varepsilon}(\theta)$ satisfying
@@
$\hat{\nabla}_{\varepsilon}F_{i}(\theta)=\frac{n}{\xi}(f_{i}(x_{\varepsilon}(\theta+\xi u),\theta+\xi u)-f_{i}(x_{\varepsilon}(\theta),\theta))u.$
@@
4   Apply Algorithm 2 to receive inexact evaluations $x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k})$ and $x_{\varepsilon_{k}}(\theta_{k})$ of the corresponding NEs of the low-level game (1).
5   Evaluate the inexact zeroth-order gradient approximation $\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ by (3).
@@
2   for i=1,,Ni=1,...,N in parallel do
3     Receive $\tilde{\theta}_{k}$ (that is, $\theta_{k}+\xi u_{k}$ or $\theta_{k}$) from the high-level regulator.
@@ 4     Update
Players send $x_{i,t_{k}},i\in[N]$ to the high-level regulator.
Algorithm 2 Distributed NE Seeking
In a nutshell, an inexact smoothing-enabled zeroth-order method is developed in Algorithm 1. At each iteration $k$, firstly, the high-level regulator selects vector $u_{k}\in\mathbb{S}^{n}$ independently and uniformly at random and then sends $\theta_{k}+\xi u_{k}$ and $\theta_{k}$ to all players. Simultaneously, the high-level regulator makes an inexact call to a distributed NE seeking algorithm, i.e., Algorithm 2, inspired by [Bianchi21] twice to compute inexact approximations of NEs $x(\theta_{k}+\xi u_{k})$ and $x(\theta_{k})$ of the low-level game (1), denoted by $x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k})$ and $x_{\varepsilon_{k}}(\theta_{k})$, respectively. Specifically, at time $t$, each player $i$ maintains variables $x_{i,t}$ and ${\bf x}_{i,t}^{j}$ to represent its own strategy and estimate of player $j$’s strategy, respectively. Define ${\bf x}_{i,t}^{i}:=x_{i,t}$ and also ${\bf x}_{j,t}^{-i}:=\operatorname*{col}\{{\bf x}_{j,t}^{l}\}_{l\in[N]\setminus i}$. Let ${\bf x}_{i,t}:=\operatorname*{col}\{{\bf x}_{i,t}^{j}\}_{j\in[N]}\in\mathbb{R}^{d}$ and ${\bf x}_{t}:=\operatorname*{col}\{{\bf x}_{i}\}_{i\in[N]}\in\mathbb{R}^{Nd}$. In the partial-decision information setting, players update their estimates based on the information received from their neighbors by applying dynamic consensus technique (21a), and subsequently adjust their own decisions through a gradient descent step (21b). Herein, it is mentioned that the partial gradient $\nabla_{i}f_{i}$ is evaluated on the local estimate $\hat{{\bf x}}_{i,t}$, while not employing the actual decision $x_{i,t}$. After receiving the estimated feedbacks $x_{\varepsilon_{k}}(\tilde{\theta_{k}}):=\operatorname*{col}\{x_{i,t_{k}}\}_{i\in[N]}$ from the players, where $\tilde{\theta}_{k}\in\{\theta_{k},\theta_{k}+\xi u_{k}\}$, the inexact zeroth-order gradient $\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ is computed as
$\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ $=\sum_{i=1}^{N}\hat{\nabla}_{\varepsilon_{k}}F_{i}(\theta_{k})$
$=\sum_{i=1}^{N}\frac{n}{\xi}(f_{i}(x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k}),\theta_{k}+\xi u_{k})$
$\quad-f_{i}(x_{\varepsilon_{k}}(\theta_{k}),\theta_{k}))u_{k}.$ (22)
At last, the high-level regulator updates its decision using gradient descent
where $\alpha>0$ is the step size to be determined.
@@
Utilizing inexact estimation of the NE of the low-level game inherently introduces a bias in the approximation of the zeroth-order gradient. This bias, if uncontrolled, has the potential to undesirably propagate and impact the overall performance of the algorithm. Notably, the degree of inexactness in $x_{\varepsilon_{k}}(\tilde{\theta}_{k})$ plays a pivotal role in the convergence analysis, which is precisely controlled by the termination criterion $t_{k}$ specified in Algorithm 2. This criterion will be formally derived in the subsequent analysis presented in Theorem 1 to establish the convergence rate.
Remark 5.
It is worthwhile noting that the regulator in the high-level does not need to acquire all information on the players. In fact, at moment $k+1$, only two function values $f_{i}(x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k}),\theta_{k}+\xi u_{k})$ and $f_{i}(x_{\varepsilon_{k}}(\theta_{k}),\theta_{k})$ are required to be disclosed by the high-level regulator to update its decision $\theta_{k}$. Moreover, compared to gradient-based algorithms [Maljkovic2024OnDC, linon24, jo2023], the regulator does not need to compute players’ gradient information, making this more aligned with real-world scenarios. In addition, it is noteworthy that in the setting of the partial-decision information regarding the underlying game, each individual strategy and gradient of the objective function are not accessible to other players. Consequently, every player must exert extra effort to estimate strategies of other players and hence steps (21a) and (21c) are introduced in Algorithm 2. However, during the practical implementation of Algorithm 1, it is only necessary to employ simple gradient descent method to update the players’ decision variables to obtain approximate solution $x_{\varepsilon_{k}}(\tilde{\theta}_{k})$ of the low-level game, which will simplify Algorithm 1 greatly.
@@
It follows from Lemma 1 in [Bianchi21] that $\mathbf{G}({\bf x},\theta)$ is Lipschitz continuous with respect to $\bf x$}, as stated formally below.
@@
Under Assumptions 1-iii) and 2-i), for any $\theta\in\mathbb{R}^{n}$, the mapping $\mathbf{G}$ is $l^{\prime}$-Lipschitz continuous for some $\mu\leq l^{\prime}\leq l$, that is,
@@
To proceed further analysis, the convergence result of Algorithm 2 for resolving the low-level game (1) is presented.
@@
with $a:=\bar{\sigma}^{2}(2ll^{\prime}+l^{\prime 2}+4\mu l^{\prime}+2l^{2})+2(l^{2}l^{\prime 2}+\mu l^{\prime 2}+2l^{2}l^{\prime 2})\bar{\sigma}^{2}+2l^{2}l^{\prime 2}\bar{\sigma}^{2}$, then
@@
Remark 6.
As is shown in (3), there exists a tradeoff between the inexactness of Algorithm 2 and the iteration number $t_{k}$, which will be carefully specified as discussed later. Besides, different from Theorem 1 in [Bianchi21], an explicit upper bound on the require stepsize $\gamma$ for eatablishing linear convergence of Algorithm 2 is given here in (24).
The following lemma reveals the Lipschitz continuous of the upper-level function $f_{i}$, which is rather crucial in analyzing the algorithm.
Lemma 5.
Under Assumptions 1 and 2, $f_{i}(x(\theta),\theta)$ is Lipschitz continuous, i.e., there exists a constant $L_{F}>0$ such that
Proof.
See Appendix C. ∎
Define a natural filtration
$\mathscr{F}_{0}:=\{\emptyset,\Omega\},$
$\mathscr{F}_{k}:=\sigma\{\theta_{0},u_{t}:0\leq t\leq k-1\},\ \forall k\geq 1.$
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]$
$\leq\frac{4(\hat{\mathbf{F}}(\theta_{0})-\hat{\mathbf{F}}^{*})}{\alpha K}+8N^{2}n^{2}L_{F}^{2}\alpha\frac{nNL_{F}+1}{\xi}$